# Operad of Posets 101: The Wixárika Posets

### v2

@@
Abstract.
We study sets whose combinatorics are related to the combinatorics of posets.
The language of operads provides us with tools to better understand the combinatorics of these objects. In this note we describe a non-trivial example of a suboperad, called the Wixárika posets, together with its associated algebras. This example is rich enough to showcase the particularities of the field, without delving into technicalities.
1. Introduction
@@
Let us begin by providing an overview of what this note is about. Let $P$ be a partially ordered set (poset). We associate to $P$ a directed acyclic simple graph, called the Hasse diagram $H_{P}$ of $P$: Each vertex of $H_{P}$ represents an element of $P$ and there is an arrow from $x$ to $y$ if $x<y$ and there is no $z$ with $x<z<y$. We can visualize compositions of posets $P$ and $Q$ in terms of their Hasse diagrams. Given a point $p\in P$, the composition $P\circ_{p}Q$ is the poset whose Hasse diagram looks like the Hasse diagram of $P$ except that the vertex $p$ is replaced by the diagram $H_{Q}$. Using this composition we can associate to the poset $\{x,y\}$, the disjoint union of two elements, a map of posets that sends $P,Q$ to $\{x,y\}(P,Q)=(\{x,y\}\circ_{x}P)\circ_{y}Q=P\sqcup Q$, the disjoint union of $P$ and $Q$.
To formalize the algebraic structure highlighted in the previous example we use the language of operads. An operad consists of sets $O(n)$ for all $n\geq 1$, and composition maps $O(n)\times O(k_{1})\times\cdots\times O(k_{n})\mapsto O(\sum_{i=1}^{n}k_{i})$, satisfying certain compatibility conditions (see Section 4). In the previous example, the operad of finite posets is formed by $\mathrm{Poset}(n)$, the sets of posets with $n$ elements. For $P\in\mathrm{Poset}(n),Q_{i}\in\mathrm{Poset}(k_{i}),i=1,\cdots,n,$ the composition $P\circ(Q_{1},\cdots,Q_{n})$ is the poset whose Hasse diagram is obtained by replacing for each $i=1,\ldots,n$ the $i$-th vertex of $H_{P}$ by $H_{Q_{i}}$.
Motivated by this construction we aim to study sets $X$ together with maps $\mathrm{Poset}(n)\mapsto\mathrm{Map}(X^{n},X)$. For a correspondence that sends a poset $P$ to a map $f_{P}$, we also require that the poset $P\circ_{p_{i}}Q$ is send to $f_{P}(id,\cdots,id,f_{Q},id,\cdots,id)$ where $f_{Q}$ is located on the $i$-th entry.
Here is an explicit example: R. Stanley [STA70] defined the strict $n$-order polynomials $\Omega^{\circ}(P,n)$ of $P$ as the number of strict order-preserving maps from $P$ to the ordered set $\langle n\rangle=\{1<\cdots<n\}$. By definition $\Omega^{\circ}(P\sqcup Q,n)=\Omega^{\circ}(P,n)\Omega^{\circ}(Q,n)$. We interpret the previous result as suggesting that the endomorphism associated with $\{x,y\}$ is the product of order polynomials, that is, $\{x,y\}(\Omega^{\circ}(P,n),\Omega^{\circ}(Q,n))=\Omega^{\circ}\left(\{x,y\}(P,Q),n\right)=\Omega^{\circ}(P,n)\Omega^{\circ}(Q,n)$. By abuse of notation we denote with $\{x,y\}$ an operation on polynomials and an operation on posets, since this choice clarifies the algebra structure.
Another example of a set in which posets act as endomorphisms comes from polytope theory. The order polytope of a poset $P$ is defined as
@@
One can verify that $\mathrm{Poly}(\{x,y\}(P,Q))=\{x,y\}(\mathrm{Poly}(P),\mathrm{Poly}(Q))$ if $\{x,y\}$ evaluated on polytopes is the Minkowski sum of polytopes.
@@
As this is an introductory text, it will be more convenient to work with a particular family of posets which we call Wixárica posets. We will then describe how topological information of these posets is transferred to the algebraic level.
The remainder is organized as follows. In Section 2 we describe posets and introduce Wixárika posets. In Section 3 we introduce an algebra over the operad of Wixárika posets: the order series. Section 4 formalizes the notion of operads and their algebras. Section 5 provides a worked-out example in enumerative combinatorics. Section 6 provides an example of the use of topological information that is transferred among algebras over an operad. We then suggest future reading in Section 7.
2. Wixárica Posets
In this section, we will define our main object of study, which is the class of Wixárica posets. We are interested in some algebraic properties of these posets. In particular, we will define the operations of concatenation and disjoint union of posets. Moreover, we will introduce a new operation, called the ’itsari operation, to obtain Wixárica posets through the consecutive application of these operations. We will look at the Hasse diagrams of Wixárica posets in order to have graphical representations.
When dealing with posets, we have two definitions: the strict and the non-strict orders. Most of the time, we are working with the second. We shall give the two for completeness.
Definition 1.
@@ 2. Wixárica Posets
Definition 3.
A set $P$ together with a (strict) partial order $(P,\leq)$ is called a (strict) partially ordered set or simply a (strict) poset. We use the notation for the set and the poset with underlying set $P$. To avoid confusion, we use $\leq_{P}$ to refer to implicit order.
Example 4.
Examples of posets. Ordered sets, such as the real numbers; the set of subsets of a given set, its power set, ordered by inclusion; the set of natural numbers equipped with the divisibility relation; if we consider a finite set of programs with the order relations $A<B$ whenever the output of $A$ is an input of $B$, then a set of programs forms a poset where the order gives restrictions to select at what turn to run each program.
In this work we assume every poset is finite.
A Hasse diagram is a graph used to represent a finite partially ordered set, in the form of a drawing of its transitive reduction. We will draw Hasse diagrams when we need to visualize certain posets. Concretely, a partially ordered set $P$ with order $\leq$ is represented by a graph whose vertices are the elements of $P$, and we have an edge from one vertex $r$ to another vertex $s$ whenever $r\neq s$ and $r<s$ and there is no $t$ with $r<t<s$. Typically, in Hasse diagrams, the greater element is positioned higher than the lesser.
Example 5.
An example is given in Figure 1, it shows the poset given by the divisors of 60.
Figure 1. Hasse diagram of the poset of divisors of 60 (Image created by Ed_g2s, https://commons.wikimedia.org/wiki/User:Ed_g2s).
@@
  1. 1)
    A single element is series-parallel.
  2. 2)
@@
Example 9.
An example is illustrated in Figure 2.
Figure 2. Hasse diagram of a series-parallel poset (Image created by David Eppstein, https://commons.wikimedia.org/wiki/User:David_Eppstein/Gallery).
@@
’Itsari is a word from Wixárica language whose meaning is to weave. Geometrically, we are adding a new minimum $x_{0}$ and a new maximum $x_{1}$ to $X$, then we attach a handle on these points, that is, we add a third element $y$ with $x_{0}<y<x_{1}$.
Definition 11.
@@
Example 12.
An example of a Wixárica poset is shown by the Hasse diagram in Figure 3. Note that in contrast to the previous examples, we direct Hasse diagrams of Wixárica posets from left to right.
Figure 3. Hasse diagram of a Wixárika poset.
@@
3. Order Series
In this section, we review the theory of power series and describe a set of power series associated to posets, as well as operations between them. The idea is that the operations introduced in this section are related to the poset operations defined in the previous section. Most of our examples of posets and ordered sets are given using a particular type of very simple poset, called a chain.
Definition 13.
Define a $n$-chain to be the total ordered set $\langle n\rangle=\{1<2<\cdots<n\}$ and let $P$ be a finite partially ordered set. An order-preserving map $\sigma:P\to\langle n\rangle$ is a map such that if $r\leq s$ in $P$, then $\sigma(r)\leq\sigma(s)$. Define $\Omega^{\circ}(P,n)$ as the number of order-preserving maps from $P$ to $\langle n\rangle$. A strictly order-preserving map $\tau:P\to\langle n\rangle$ is a map such that if $r<s$ then $\tau(r)<\tau(s)$. Define $\Omega(P,n)$ to be the number of strictly order-preserving maps from $P$ to $\langle n\rangle$.
Definition 14.
[…]
Example 15.
[…]
Definition 16.
For a poset $P$ we define a formal power series, called the order series of $P$, by
@@ 3. Order Series
(1) $\mathfrak{Z}\left(m\right):=\mathfrak{Z}\left(\langle m\rangle\right)=\sum_{n=m}^{\infty}{n\choose m}x^{n}=\frac{x^{m}}{(1-x)^{m+1}}.$
For a proof of the second equality see [WIL06, Equation (1.5.5)], [GOU72, Equation (1.3)] or [PET15, Equation (1.3)].
@@
$\left(\sum_{i=0}^{\infty}a_{i}x^{i}\right)\left(\sum_{i=0}^{\infty}b_{i}x^{i}\right)=\sum_{n=0}^{\infty}c_{n}x^{n}\quad\text{where}\quad c_{n}=\sum_{k==0}^{n}a_{k}b_{n-k}.$
Definition 19.
If $x$ is the common variable for the order series $\mathfrak{Z}\left(Q\right)$ and $\mathfrak{Z}\left(P\right)$, we define the direct product using the Cauchy product as follows:
(2) $\mathfrak{Z}\left(Q\right)*\mathfrak{Z}\left(P\right)=\mathfrak{Z}\left(Q\right)(1-x)\mathfrak{Z}\left(P\right).$
Definition 20.
@@
In the case of order series we will call the Hadamard product disjoint union and use the symbol $\sqcup$, that is,
$\mathfrak{Z}\left(P\right)\sqcup\mathfrak{Z}\left(Q\right)=\mathfrak{Z}\left(P\right)\odot\mathfrak{Z}\left(Q\right).$
Definition 21.
@@
$D(\mathfrak{Z}\left(P\right))=\mathfrak{Z}\left(D(P)\right).$
For the case of chains $\langle n\rangle$ we denote $D(\langle n\rangle)=D(n)$.
Note that we are abusing notation: We are using the same symbol for operations between posets and order series. The meaning, however, will be clear from the context.
The following two results are well-known (see [FS86], [WAG92], [BRA06], [BHK18],[ABD23]).
Proposition 22.
@@
$\mathfrak{Z}\left(P*Q\right)=\mathfrak{Z}\left(P\right)(1-x)\mathfrak{Z}\left(Q\right).$
Proposition 23.
@@
Proposition 25.
Unraveling the definitions, we can show linearity of $D$:
$D(a\mathfrak{Z}\left(P\right)+b\mathfrak{Z}\left(Q\right))=aD(\mathfrak{Z}\left(P\right))+bD(\mathfrak{Z}\left(Q\right)).$
Example 26.
@@
$\mathfrak{Z}\left(D({\langle n\rangle})\right)$ $=\mathfrak{Z}\left(1\right)*\mathfrak{Z}\left(\langle 1\rangle\sqcup\langle n\rangle\right)*\mathfrak{Z}\left(1\right)$
$=\frac{x}{(1-x)^{2}}*(n\frac{x^{n}}{(1-x)^{n+1}}+(n+1)\frac{x^{n+1}}{(1-x)^{n+2}})*\frac{x}{(1-x)^{2}}$
@@
4. Algebras over the operad of Wixárika posets
In this section, we introduce the language of operads. This allows us to compare sets in which the same operations are defined, similarly to homomorphisms of groups.
Definition 27.

### v3

@@
Abstract
We study classes of objects whose combinatorics are closely related to those of posets. The framework of operads and operad algebras allows us to make this relationship precise and provides tools for a deeper understanding of their combinatorial structure.
In this note, we present a nontrivial example of a suboperad of the operad of posets, called Wixárika posets, together with its associated algebras. This example is sufficiently rich to exhibit key structural features of the theory, while remaining accessible and avoiding unnecessary technicalities.
1 Introduction
@@
Let us begin with an overview of what this note is about. Let $P$ be a partially ordered set (poset). We associate to $P$ a directed acyclic simple graph, the Hasse diagram $H_{P}$ of $P$: Each vertex of $H_{P}$ represents an element of $P$ and there is an arrow from $x$ to $y$ if $x<y$ and there is no $z$ with $x<z<y$. We can visualize the composition of posets $P$ and $Q$ in terms of their Hasse diagrams. Given a point $p\in P$, the composition $P\circ_{p}Q$ is the poset whose Hasse diagram is the Hasse diagram of $P$ with the vertex $p$ replaced by the diagram $H_{Q}$.
Using this composition, we can associate to the poset $\{x,y\}$, the disjoint union of two elements, a map of posets that sends $P,Q$ to their disjoint union,
$\{x,y\}_{\hbox{Poset}}(P,Q)=(\{x,y\}_{\hbox{Poset}}\circ_{x}P)\circ_{y}Q=P\sqcup Q.$
This algebraic structure can be formalized using the language of operads. An operad consists of sets $O(n)$ for all $n\geq 1$, together with composition maps $O(n)\times O(k_{1})\times\cdots\times O(k_{n})\mapsto O(\sum_{i=1}^{n}k_{i})$, satisfying certain compatibility conditions (see Section 4). In the previous example, the operad of finite posets is formed by $\mathrm{Poset}(n)$, the sets of posets with $n$ elements. For $P\in\mathrm{Poset}(n),Q_{i}\in\mathrm{Poset}(k_{i}),i=1,\cdots,n,$ the composition $P\circ(Q_{1},\cdots,Q_{n})$ is the poset whose Hasse diagram is obtained by replacing for each $i=1,\ldots,n$ the $i$-th vertex of $H_{P}$ by $H_{Q_{i}}$.
Thus, from the operadic point of view each poset also represents an endomorphism of posets. Motivated by this construction, we aim to study sets $X$ together with maps $\mathrm{Poset}(n)\mapsto\mathrm{Map}(X^{n},X)$. For a correspondence that sends a poset $P$ to a map $f_{P}$, we also require that the poset $P\circ_{p_{i}}Q$ is sent to the map composition $f_{P}(id,\cdots,id,f_{Q},id,\cdots,id)$ where $f_{Q}$ is located on the $i$-th entry.
Here is an explicit example: R. Stanley [22] defined the strict order polynomial $\Omega^{\circ}(P,n)$ of $P$ by counting the number of strict order-preserving maps from $P$ to the ordered sets $\langle n\rangle=\{1<\cdots<n\},n\in\mathbb{N}$. By definition, $\Omega^{\circ}(P\sqcup Q,n)=\Omega^{\circ}(P,n)\Omega^{\circ}(Q,n)$. We rephrase this result by saying that the endomorphism associated with $\{x,y\}$ is the product of strict order polynomials, that is,
$\{x,y\}_{\hbox{Polynomials}}(\Omega^{\circ}(P,n),\Omega^{\circ}(Q,n))=\Omega^{\circ}\left(\{x,y\}_{\hbox{Poset}}(P,Q),n\right)=\Omega^{\circ}(P,n)\Omega^{\circ}(Q,n).$
Note that here and in what follows we include a subscript to indicate the set on which a poset acts.
Another example comes from polytope theory. The order polytope of a poset $P$ is defined as
@@
One can verify that
$\mathrm{Poly}(\{x,y\}_{\hbox{Poset}}(P,Q))$ $=\{x,y\}_{\hbox{Polytope}}(\mathrm{Poly}(P),\mathrm{Poly}(Q))$
$=\{v+w\mid v\in\mathrm{Poly}(P),w\in\mathrm{Poly}(Q)\},$
that is, the map $\{x,y\}_{\hbox{Polytope}}$ is the Minkowski sum of polytopes.
@@
As this is an introductory text, it will be more convenient to work with a particular family of posets which we call Wixárica posets. We outline the theory for this case. As an application, we describe how topological information of these posets (i.e., of their Hasse diagrams) is reflected at the algebraic level.
The exposition is organized as follows. In Section 2 we define posets and introduce Wixárika posets. In Section 3 we introduce an algebra over the operad of Wixárika posets: Generating functions of the above mentioned order polynomials, called order series. Section 4 formalizes the notion of operads and their algebras. Section 5 provides a worked-out example in enumerative combinatorics. Section 6 gives an example of how topological information is transferred from the Wixárika operad to one of its algebras. We then suggest future reading in Section 7.
2 Wixárica Posets
In this section, we define our main object of study, the class of Wixárica posets. We are interested in algebraic properties of these posets. In particular, we will define the operations of concatenation and disjoint union of posets. Moreover, we will introduce a new operation, called the ’itsari operation, to obtain Wixárica posets through the consecutive application of the latter operations. We then interpret these in terms of Hasse diagrams of Wixárica posets in order to obtain graphical representations of these operations.
In principle, posets may be defined using two variants of order relations, either strict or non-strict orders. Most of the time, we work with the latter. We shall give both definitions for completeness.
Definition 2.1.
@@ 2 Wixárica Posets
Definition 2.3.
A set $P$ together with a (strict) partial order $(P,\leq)$ is called a (strict) partially ordered set or simply a (strict) poset. We use $P$ to denote both the underlying set and the poset. If confusion is possible, we add a subscript (e.g., $\leq_{P}$) to distinguish different partial orders.
Example 2.4.
Examples of posets: Ordered sets, such as the real numbers; the set of subsets of a given set, its power set, ordered by inclusion; the set of natural numbers equipped with the divisibility relation; if we consider a finite set of programs with the order relations $A<B$ whenever the output of $A$ is an input of $B$, then a set of programs forms a poset where the order gives restrictions to select at what turn to run each program.
In this note we assume that every poset is finite.
A Hasse diagram is a graph that represents a finite poset by drawing its transitive reduction. We will draw Hasse diagrams when we need to visualize certain posets. Concretely, a partially ordered set $P$ with order $\leq$ is represented by a graph whose vertices are the elements of $P$, and we have an edge from one vertex $r$ to another vertex $s$ whenever $r\neq s$ and $r<s$ and there is no $t$ with $r<t<s$. Typically, in Hasse diagrams the greater element is positioned higher than the lesser.
Example 2.5.
An example is given in Figure 1. It shows the poset given by the divisors of 60. There is an edge from 1 to the numbers 2, 3, and 5, since 1 is their closest divisor. For the remaining numbers, we can apply transitivity. This means that 1 is a divisor of 5, and 5 is a divisor of 15, but an edge between 1 and 15 is not required. Note also that 1 is located at the bottom, since it is the smallest number. Each number is placed on a higher level according to its order. For this reason, 60 is located at the top.
Figure 1: Hasse diagram of the poset of all divisors of 60. Edges denotes the order relation, and higher position encodes the hierarchy (Image created by Ed_g2s, https://commons.wikimedia.org/wiki/User:Ed_g2s).
@@
Recall from the introduction how the disjoint union $P\sqcup Q$ can be formulated using poset composition and the additional poset $\{x,y\}$, a disjoint union of two elements:
$P\sqcup Q=\{x,y\}_{\hbox{Poset}}(P,Q)=(\{x,y\}_{\hbox{Poset}}\circ_{x}P)\circ_{y}Q.$
In the same way the poset $\{x<y\}$ gives rise to the direct sum,
$P*Q=\{x<y\}_{\hbox{Poset}}(P,Q)=(\{x<y\}_{\hbox{Poset}}\circ_{x}P)\circ_{y}Q.$
From now on we stick to the symbols $\sqcup$ and * to keep the notation light. Nevertheless, it is instructive in what follows to keep in mind that these operations on posets are induced by posets themselves.
@@
  1. 1.
    A single point is series-parallel.
  2. 2.
@@
Example 2.9.
An example is illustrated in Figure 2. The central blue rectangle illustrates parallel composition, that is, a disjoint union of three posets, one in each brown rectangle. The final poset is obtained by two series compositions, where in each case edges are added from the maximal element(s) of one blue rectangle to the minima element(s) of the next blue rectangle.
Figure 2: Hasse diagram of a series-parallel poset. Posets in brown rectangles are joined together by disjoint union (parallel composition), posets in blue rectangles are joined by direct sum (series composition). (Image created by David Eppstein, https://commons.wikimedia.org/wiki/User:David_Eppstein/Gallery).
@@
In the case of chains $\langle n\rangle$ we abbreviate $D(\langle n\rangle)$ by $D(n)$ (cf. Section 5).
’Itsari is a word from Wixárica language that means to weave. Geometrically, we are adding a new minimum $x_{0}$ and a new maximum $x_{1}$ to $X$, then we attach a handle on these points, that is, we add a third element $y$ with $x_{0}<y<x_{1}$.
Definition 2.11.
@@
Example 2.12.
An example of a Wixárica poset is shown by the Hasse diagram in Figure 3. We start by building a chain using the direct sum. Then we add arcs using the ’itsari operation. Note that in contrast to the previous examples, we direct Hasse diagrams of Wixárica posets from left to right.
Figure 3: Hasse diagram of a Wixárika poset, oriented from left to right.
@@
3 Order Series
In this section, we describe a set of power series associated to posets and operations between them. The idea is that these operations are related to the poset operations defined in the previous section.
Most of our examples of posets and ordered sets use a simple type of poset, called a chain.
Definition 3.1.
Define a $n$-chain to be the total ordered set $\langle n\rangle=\{1<2<\cdots<n\}$. Let $P$ be a finite partially ordered set. An order-preserving map $\sigma:P\to\langle n\rangle$ is a map such that if $r\leq s$ in $P$, then $\sigma(r)\leq\sigma(s)$. Define $\Omega(P,n)$ as the number of order-preserving maps from $P$ to $\langle n\rangle$. A strictly order-preserving map $\tau:P\to\langle n\rangle$ is a map such that if $r<s$ then $\tau(r)<\tau(s)$. Define $\Omega^{\circ}(P,n)$ to be the number of strictly order-preserving maps from $P$ to $\langle n\rangle$.
Definition 3.2.
Given $n,m\in\mathbb{N}$ define the multiset $\left(\kern-3.00003pt\binom{n}{m}\kern-3.00003pt\right):={n+m-1\choose m}$.
Example 3.3.
If $P=\langle m\rangle$ we have $\Omega^{\circ}(\langle m\rangle,n)={n\choose m}$, while $\Omega(\langle m\rangle,n)=\left(\kern-3.00003pt\binom{n}{m}\kern-3.00003pt\right)$. If $P$ is the disjoint union of $p$ points, then $\Omega^{\circ}(P,n)=\Omega(P,n)=n^{p}$.
Definition 3.4.
For a poset $P$ we define a formal power series, called the (strict) order series of $P$, by
@@ 3 Order Series
$\mathfrak{Z}\left(m\right):=\mathfrak{Z}\left(\langle m\rangle\right)=\sum_{n=m}^{\infty}{n\choose m}x^{n}=\frac{x^{m}}{(1-x)^{m+1}}.$ (3.1)
For a proof of the second equality see [27, Equation (1.5.5)], [12, Equation (1.3)] or [20, Equation (1.3)].
@@
$\left(\sum_{i=0}^{\infty}a_{i}x^{i}\right)\left(\sum_{i=0}^{\infty}b_{i}x^{i}\right)=\sum_{n=0}^{\infty}c_{n}x^{n}\quad\text{where}\quad c_{n}=\sum_{k=0}^{n}a_{k}b_{n-k}.$
Definition 3.7.
If $x$ is the common variable for the order series $\mathfrak{Z}\left(P\right)$ and $\mathfrak{Z}\left(Q\right)$, we define the direct product using the Cauchy product as follows:
$\mathfrak{Z}\left(P\right)\hat{*}\mathfrak{Z}\left(Q\right)=\mathfrak{Z}\left(P\right)(1-x)\mathfrak{Z}\left(Q\right).$ (3.2)
Definition 3.8.
@@
[…]
Definition 3.9.
@@
$\widehat{D}(\mathfrak{Z}\left(P\right))=\mathfrak{Z}\left(D(P)\right).$
In the case of chains $\langle n\rangle$ we abbreviate $\widehat{D}(\mathfrak{Z}\left(n\right))$ by $\widehat{D}(n)$ (cf. Section 5).
The following two formulas are well-known (see [9], [26], [5], [6],[1]).
Proposition 3.10.
@@
$\mathfrak{Z}\left(P*Q\right)=\mathfrak{Z}\left(P\right)\hat{*}\mathfrak{Z}\left(Q\right)=\mathfrak{Z}\left(P\right)(1-x)\mathfrak{Z}\left(Q\right),$ (3.3)
[…]
Note that Equation (3.4) follows directly from counting the number of order preserving maps from $P\sqcup Q$ to $n$. In the case of chains, the paper [1] provides an explicit formula for the order series of disjoint unions:
Proposition 3.11.
@@
Proposition 3.13.
Unraveling the definitions, we can show linearity of $\widehat{D}$:
$\widehat{D}(a\mathfrak{Z}\left(P\right)+b\mathfrak{Z}\left(Q\right))=a\widehat{D}(\mathfrak{Z}\left(P\right))+b\widehat{D}(\mathfrak{Z}\left(Q\right)).$
Example 3.14.
@@
$\mathfrak{Z}\left(D(n)\right)$ $=\mathfrak{Z}\left(1\right)\hat{*}\mathfrak{Z}\left(\langle 1\rangle\sqcup\langle n\rangle\right)\hat{*}\mathfrak{Z}\left(1\right)$
$=\frac{x}{(1-x)^{2}}\hat{*}\left(n\frac{x^{n}}{(1-x)^{n+1}}+(n+1)\frac{x^{n+1}}{(1-x)^{n+2}}\right)\hat{*}\frac{x}{(1-x)^{2}}$
@@
4 Algebras over the operad of Wixárika posets
In this section, we introduce the language of operads. This allows us to compare sets in which the same operations are defined, similarly to groups and group homomorphisms.
Definition 4.1.